# Local-Cloud Inference Offloading for LLMs in Multi-Modal, Multi-Task, Multi-Dialogue Settings

### v3

@@ 4.3. Main Experimental Results
Comparison with SOTA Exploration-Decision Baselines. Existing works do not consider multi-task and multi-dialogue settings in a multi-modal scenario, and cannot be trivially deployed to our setting. For example, (He et al., 2024) utilize a benchmark with a fixed evaluation metric Pass@k that measures success rates. In contrast, our method addresses a multi-modal multi-task and multi-dialogue setting which introduces significant challenges related to response score evaluation and uncertainty. From the results, our TMO system significantly outperforms these SOTA baselines. (He et al., 2024) fails to efficiently select LLMs and modalities, frequently opting to upload all three types of modality data sources for most actions. This results in substantial resource waste and redundancy in uploading unnecessary information to the cloud. Conversely, (Yang et al., 2024) rarely considers any modality data, relying solely on the local LLM and text-only cloud LLM. Although this approach avoids violating any resource constraints, it leads to lower response scores, as the LLM lacks access to any information derived from modality data.
@@ Acknowledgements.
This work was supported in part by the National Science Foundation (NSF) under grant CPS-2313109, by the Office of Naval Research (ONR) under grant N00014-23-C-1016, and by the Air Force Office of Scientific Research (AFOSR) under grant FA9550-24-1-0083.

### v4

@@ 4.3. Main Experimental Results
Comparison with SOTA Exploration-Decision Baselines. Existing works do not consider multi-task and multi-dialogue settings in a multi-modal scenario, and cannot be trivially deployed to our setting. For example, (he2024large) utilize a benchmark with a fixed evaluation metric Pass@k that measures success rates. In contrast, our method addresses a multi-modal multi-task and multi-dialogue setting which introduces significant challenges related to response score evaluation and uncertainty. From the results, our TMO system significantly outperforms these SOTA baselines. (he2024large) fails to efficiently select LLMs and modalities, frequently opting to upload all three types of modality data sources for most actions. This results in substantial resource waste and redundancy in uploading unnecessary information to the cloud. Conversely, (yang2024perllm) rarely considers any modality data, relying solely on the local LLM and text-only cloud LLM. Although this approach avoids violating any resource constraints, it leads to lower response scores, as the LLM lacks access to any information derived from modality data.
@@ Acknowledgements.
This work was supported in part by the National Science Foundation (NSF) under grant CPS-2313109, by the Office of Naval Research (ONR) under grant N00014-23-C-1016, by the Air Force Office of Scientific Research (AFOSR) under grant FA9550-24-1-0083, and by NVIDIA’s Academic Grant Program.